OpenStudy (anonymous):

Solve log11 (y + 8) + log11 4 = log11 60

OpenStudy (anonymous):

Solve log11 (y + 8) + log11 4 = log11 60 y = 7 y = 13 y = 28 y = 48

OpenStudy (raden):

use the property of logarithm : log_m a + log_m b = log_m (ab)

OpenStudy (raden):

so, log11 (y + 8) + log11 4 = log11 60 it can rewriten be log11 (y + 8)4 = log11 60 now, cancel out the log 11 to both sides, giving us (y+8)4 = 60 solve for y

OpenStudy (anonymous):

so it would equal 7?

OpenStudy (raden):

exactly, yes!

OpenStudy (anonymous):

thank you

OpenStudy (raden):

you're welcome